# Binding Free Energies without Alchemy

### 12 Mar 2026 v1

@@ 1 Introduction
On the other hand, absolute binding affinity (ABFE) calculations utilize rigorous statistical mechanics to predict the Gibbs free energy of binding. Several ABFE methods have been developed, most notably double decoupling (DD) 25. All such methods are computationally expensive due to the need to run many molecular dynamics simulations on “alchemically modified” states; however, they are highly correlated with experimental binding affinities. Thus, ABFE has emerged as the gold standard computational technique to predict protein-ligand binding affinities in the absence of experimental data about the protein target 3, 22, 2. While ABFE is far too slow to screen millions of compounds directly, it can be used to re-rank the top-scoring hits from an initial screen 18, 10. MD simulations for ABFE typically employ explicit waters, though ABFE with implicit solvent has been suggested as a faster alternative 42.
@@
In this work, we propose Direct Binding Free Energy (DBFE), a new implicit solvent ABFE method that utilizes only end-state simulation data; no alchemical intermediates are needed. DBFE requires only three simulations: the receptor-only, ligand-only, and receptor-ligand complex simulations. Since the receptor and ligand simulations can be precomputed and cached, the per-ligand cost reduces to a single complex simulation compared to the many complex lambda windows needed for DD. We evaluate DBFE on both host-guest and protein-ligand benchmarks, finding that it outperforms OBC2 DD on host-guest systems in correlation and performs comparably to OBC2 MM/GBSA on protein-ligand systems. Our results suggest that the primary accuracy bottleneck for implicit solvent methods on protein-ligand systems is the solvent model itself, rather than conformational entropy.
@@ 2.2 The Direct Binding Free Energy method
We desire to compute the Gibbs free energy $\Delta G$ for the transformation $P+L\rightarrow PL$, where $P$ is a protein, $L$ is a small molecule ligand, and $PL$ is their complex. We can do this by computing the free energy difference between a system with a “decoupled” potential energy $U_{0}(x)=U_{P}(x_{P})+U_{L}(x_{L})$ and a system with the coupled potential energy $U_{1}(x)=U_{PL}(x)$. We assume we have run MD simulations of the protein, ligand, and protein-ligand complex; thus we have access to samples from $U_{P}$, $U_{L}$, and $U_{PL}$.
@@
The thermodynamic cycle we use for this method is shown in Figure 1. In the first leg of the cycle, we seek to find the free energy of restraining the relative translation and rotation between the protein and the ligand. In this work, we define $\zeta$ to be the rigid-body transformation between the current ligand coordinates $x_{L}$ and the ligand coordinates of an arbitrary reference frame from the $PL$ simulation, after the pocket $\alpha$ carbons of the current frame have been aligned with this same reference frame. In this work, all ligand transformations account for molecular symmetry (i.e., equivalent atom permutations) via modified code from spyrmsd 35. The restrained potential energy function is thus:
@@
Here we have separated $\zeta=[t,q]^{T}$ into a translation vector $t$ and a rotation quaternion $q$. The first term is a multivariate normal distribution over translations with mean $\mu$ and covariance $\Sigma$. The second term is a Bingham distribution over rotations 7; the Bingham distribution is the analogue of the Gaussian on the unit quaternion sphere $S^{3}$ (and thus on the rotation group SO(3)), parameterized by an orthogonal matrix $M$ defining the principal axes and a diagonal matrix $Z$ controlling the concentration along each axis. This potential enables us to easily sample from $U_{\zeta}$ and analytically compute the free energy $\Delta G_{0\rightarrow r}$. We fit the parameters $M$, $Z$, $\Sigma$, and $\mu$ from the largest mode of the $PL$ trajectory (computed via the MeanShift function from scikit-learn 23, 40).
@@ 3 Benchmarking DBFE
We evaluate DBFE on two benchmarks: a host-guest benchmark of cyclodextrin and octa-acid systems, and a protein-ligand benchmark of drug-like complexes. Validating free energy methods on host-guest benchmarks is a standard practice due to their simplicity, while protein-ligand benchmarks test the ability of the algorithm to scale to real-world tasks. Both the benchmarks we used for this study were also used by Setiadi et al. 42 in their work on optimizing force-field parameters for implicit solvent free energy calculations. For both benchmarks, we compare DBFE against Onufriev-Bashford-Case model II (OBC2) 38 double decoupling (OBC2 DD), TIP3P 28 DD, and OBC2 MM/GBSA (computed from the same end-state simulations as DBFE, but without conformational entropy corrections).
To perform the OBC2 DD calculations, we used Yank 14, which utilizes Hamiltonian replica exchange simulations 47 and uses MBAR to compute the final $\Delta G$ 43. We used ParmED 39 to remove explicit waters and add the OBC2 solvation model 38. Each Yank lambda window was run for 2 ns at 300K with a 2 fs Langevin integrator.
@@ 3.1 Host-Guest Benchmark
The OBC2 DD calculations used 26 complex lambda windows and 8 solvent lambda windows (34 total) per complex. In contrast, DBFE required only 3 end-state simulations (host, guest, and host-guest complex).
@@
The results are shown in Table 1 and Figure 2. DBFE achieves a Pearson correlation of rr=0.58 with experiment, outperforming OBC2 DD (rr=0.48) on these systems. OBC2 MM/GBSA performs poorly (rr=0.33), suggesting that the conformational entropy correction provided by DBFE is important for these host-guest systems. As expected, TIP3P DD performs substantially better (rr=0.89), highlighting the importance of explicit water treatment even for these relatively simple systems.
@@ 3.2 Protein-Ligand Benchmark
Due to steric clashes in the initial simulation geometry, we removed three complexes from the analysis (CyclophilinD:27, MCL-1:15, and PWWP1:15). We additionally excluded two complexes (MCL-1:60 and MCL-1:65) where the protein binding site closed during the protein-only simulation, preventing DBFE from finding any samples that passed the Lennard-Jones cutoff. We include previous results from Alibay et al. 3 for TIP3P DD calculations.
The OBC2 DD calculations used 26 complex lambda windows and 4 solvent lambda windows (30 total) per complex.
@@
We succeeded in running DBFE on 54 of the 56 complexes. The results are shown in Table 2 and Figure 3. On this benchmark, DBFE achieved rr=0.65 and $\rho$=0.64, which is slightly worse than OBC2 MM/GBSA (rr=0.71, $\rho$=0.70). OBC2 MM/GBSA computes the binding free energy from the same implicit solvent simulations without the conformational entropy correction that DBFE provides; that it performs slightly better suggests that the conformational entropy estimate from DBFE introduces noise on these protein-ligand systems. Both methods performed worse than OBC2 DD (rr=0.73), which in turn performed considerably worse than TIP3P DD (rr=0.88).
@@ 4 Discussion
On the host-guest benchmark, DBFE (rr=0.58) outperformed OBC2 DD (rr=0.48) in correlation with experiment, while OBC2 MM/GBSA performed substantially worse (rr=0.31). This suggests that for host-guest systems, the conformational entropy correction provided by DBFE is important and that the combinatorial sampling strategy of DBFE is effective for these systems with smaller, well-defined binding sites.
On the protein-ligand benchmark, DBFE (rr=0.65) performed slightly worse than OBC2 MM/GBSA (rr=0.71), which computes the binding free energy from the same implicit solvent simulations without conformational entropy corrections. This contrasts with the host-guest results and indicates that the conformational entropy estimate from DBFE introduces noise on these more complex protein-ligand systems. Both methods performed worse than OBC2 DD (rr=0.73), while TIP3P DD (rr=0.88) substantially outperformed all OBC2 methods. The gap between OBC2 and TIP3P DD (rr=0.73 vs rr=0.88) is much larger than the gap between DBFE and OBC2 DD (rr=0.65 vs rr=0.73), suggesting that improving implicit solvent models would yield greater accuracy gains than improving the free energy estimator.
DBFE’s primary advantage is computational efficiency. While DBFE’s total simulation time for a single complex can exceed that of DD (due to the long receptor simulation), this cost is amortized in a virtual screening context: the receptor simulation is run once and reused for all candidate ligands. The per-ligand cost of DBFE is then dominated by a single short complex simulation, compared to the 26 complex lambda windows used by OBC2 DD in our benchmark: a 26x reduction in per-ligand simulation cost. The exact speedup will vary with the DD protocol, as the number of lambda windows required depends on the system and the desired precision.

### 17 Mar 2026 v2

@@ 1 Introduction
On the other hand, absolute binding affinity (ABFE) calculations utilize rigorous statistical mechanics to predict the Gibbs free energy of binding. Several ABFE methods have been developed, most notably double decoupling ($D$) 25. All such methods are computationally expensive due to the need to run many molecular dynamics simulations on “alchemically modified” states; however, they are highly correlated with experimental binding affinities. Thus, ABFE has emerged as the gold standard computational technique to predict protein-ligand binding affinities in the absence of experimental data about the protein target 3, 22, 2. While ABFE is far too slow to screen millions of compounds directly, it can be used to re-rank the top-scoring hits from an initial screen 18, 10. MD simulations for ABFE typically employ explicit waters, though ABFE with implicit solvent has been suggested as a faster alternative 42.
@@
In this work, we propose Direct Binding Free Energy (DBFE), a new implicit solvent ABFE method that utilizes only end-state simulation data; no alchemical intermediates are needed. DBFE requires only three simulations: the receptor-only, ligand-only, and receptor-ligand complex simulations. Since the receptor and ligand simulations can be precomputed and cached, the per-ligand cost reduces to a single complex simulation compared to the many complex lambda windows needed for $D$. We evaluate DBFE on both host-guest and protein-ligand benchmarks, finding that it outperforms OBC2 $D$ on host-guest systems in correlation and performs comparably to OBC2 MM/GBSA on protein-ligand systems. Our results suggest that the primary accuracy bottleneck for implicit solvent methods on protein-ligand systems is the solvent model itself, rather than conformational entropy.
@@ 2.2 The Direct Binding Free Energy method
We desire to compute the Gibbs free energy $\Delta G$ for the transformation $P+L\rightarrow PL$, where $P$ is a protein, $L$ is a small molecule ligand, and PLPL is their complex. We can do this by computing the free energy difference between a system with a “decoupled” potential energy $U_{0}(x)=U_{P}(x_{P})+U_{L}(x_{L})$ and a system with the coupled potential energy $U_{1}(x)=U_{PL}(x)$. We assume we have run MD simulations of the protein, ligand, and protein-ligand complex; thus we have access to samples from $U_{P}$, $U_{L}$, and $U_{PL}$.
@@
The thermodynamic cycle we use for this method is shown in Figure 1. In the first leg of the cycle, we seek to find the free energy of restraining the relative translation and rotation between the protein and the ligand. In this work, we define $\zeta$ to be the rigid-body transformation between the current ligand coordinates $x_{L}$ and the ligand coordinates of an arbitrary reference frame from the PLPL simulation, after the pocket $\alpha$ carbons of the current frame have been aligned with this same reference frame. In this work, all ligand transformations account for molecular symmetry (i.e., equivalent atom permutations) via modified code from spyrmsd 35. The restrained potential energy function is thus:
@@
Here we have separated $\zeta=[t,q]^{T}$ into a translation vector $t$ and a rotation quaternion $q$. The first term is a multivariate normal distribution over translations with mean $\mu$ and covariance $\Sigma$. The second term is a Bingham distribution over rotations 7; the Bingham distribution is the analogue of the Gaussian on the unit quaternion sphere $S^{3}$ (and thus on the rotation group SO(3)), parameterized by an orthogonal matrix $M$ defining the principal axes and a diagonal matrix $Z$ controlling the concentration along each axis. This potential enables us to easily sample from $U_{\zeta}$ and analytically compute the free energy $\Delta G_{0\rightarrow r}$. We fit the parameters $M$, $Z$, $\Sigma$, and $\mu$ from the largest mode of the PLPL trajectory (computed via the MeanShift function from scikit-learn 23, 40).
@@ 3 Benchmarking DBFE
We evaluate DBFE on two benchmarks: a host-guest benchmark of cyclodextrin and octa-acid systems, and a protein-ligand benchmark of drug-like complexes. Validating free energy methods on host-guest benchmarks is a standard practice due to their simplicity, while protein-ligand benchmarks test the ability of the algorithm to scale to real-world tasks. Both the benchmarks we used for this study were also used by Setiadi et al. 42 in their work on optimizing force-field parameters for implicit solvent free energy calculations. For both benchmarks, we compare DBFE against Onufriev-Bashford-Case model II (OBC2) 38 double decoupling (OBC2 $D$), TIP3P 28 $D$, and OBC2 MM/GBSA (computed from the same end-state simulations as DBFE, but without conformational entropy corrections).
To perform the OBC2 $D$ calculations, we used Yank 14, which utilizes Hamiltonian replica exchange simulations 47 and uses MBAR to compute the final $\Delta G$ 43. We used ParmED 39 to remove explicit waters and add the OBC2 solvation model 38. Each Yank lambda window was run for 2 ns at 300K with a 2 fs Langevin integrator.
@@ 3.1 Host-Guest Benchmark
The OBC2 $D$ calculations used 26 complex lambda windows and 8 solvent lambda windows (34 total) per complex. In contrast, DBFE required only 3 end-state simulations (host, guest, and host-guest complex).
@@
The results are shown in Table 1 and Figure 2. DBFE achieves a Pearson correlation of rr=0.58 with experiment, outperforming OBC2 $D$ (rr=0.48) on these systems. OBC2 MM/GBSA performs poorly (rr=0.33), suggesting that the conformational entropy correction provided by DBFE is important for these host-guest systems. As expected, TIP3P $D$ performs substantially better (rr=0.89), highlighting the importance of explicit water treatment even for these relatively simple systems.
@@ 3.2 Protein-Ligand Benchmark
Due to steric clashes in the initial simulation geometry, we removed three complexes from the analysis (CyclophilinD:27, MCL-1:15, and PWWP1:15). We additionally excluded two complexes (MCL-1:60 and MCL-1:65) where the protein binding site closed during the protein-only simulation, preventing DBFE from finding any samples that passed the Lennard-Jones cutoff. We include previous results from Alibay et al. 3 for TIP3P $D$ calculations.
The OBC2 $D$ calculations used 26 complex lambda windows and 4 solvent lambda windows (30 total) per complex.
@@
We succeeded in running DBFE on 54 of the 56 complexes. The results are shown in Table 2 and Figure 3. On this benchmark, DBFE achieved rr=0.65 and $\rho$=0.64, which is slightly worse than OBC2 MM/GBSA (rr=0.71, $\rho$=0.70). OBC2 MM/GBSA computes the binding free energy from the same implicit solvent simulations without the conformational entropy correction that DBFE provides; that it performs slightly better suggests that the conformational entropy estimate from DBFE introduces noise on these protein-ligand systems. Both methods performed worse than OBC2 $D$ (rr=0.73), which in turn performed considerably worse than TIP3P $D$ (rr=0.88).
@@ 4 Discussion
On the host-guest benchmark, DBFE (rr=0.58) outperformed OBC2 $D$ (rr=0.48) in correlation with experiment, while OBC2 MM/GBSA performed substantially worse (rr=0.31). This suggests that for host-guest systems, the conformational entropy correction provided by DBFE is important and that the combinatorial sampling strategy of DBFE is effective for these systems with smaller, well-defined binding sites.
On the protein-ligand benchmark, DBFE (rr=0.65) performed slightly worse than OBC2 MM/GBSA (rr=0.71), which computes the binding free energy from the same implicit solvent simulations without conformational entropy corrections. This contrasts with the host-guest results and indicates that the conformational entropy estimate from DBFE introduces noise on these more complex protein-ligand systems. Both methods performed worse than OBC2 $D$ (rr=0.73), while TIP3P $D$ (rr=0.88) substantially outperformed all OBC2 methods. The gap between OBC2 and TIP3P $D$ (rr=0.73 vs rr=0.88) is much larger than the gap between DBFE and OBC2 $D$ (rr=0.65 vs rr=0.73), suggesting that improving implicit solvent models would yield greater accuracy gains than improving the free energy estimator.
DBFE’s primary advantage is computational efficiency. While DBFE’s total simulation time for a single complex can exceed that of $D$ (due to the long receptor simulation), this cost is amortized in a virtual screening context: the receptor simulation is run once and reused for all candidate ligands. The per-ligand cost of DBFE is then dominated by a single short complex simulation, compared to the 26 complex lambda windows used by OBC2 $D$ in our benchmark: a 26x reduction in per-ligand simulation cost. The exact speedup will vary with the $D$ protocol, as the number of lambda windows required depends on the system and the desired precision.